TurtlePrincess:

There are 5 seniors and 3 juniors on the debate team. What is the probability that a 3 person team chosen to go to a championship will consist of all seniors? A. 1/4,032 B. 5/28 C. 3/8

Shawnte:

there's 8 people in the group 3 of them are jinors....this is a simple fraction

TurtlePrincess:

Thank you!

Shawnte:

no problem

Shawnte:

wait do you know what the answer is.......... i basically told you the fraction

TurtlePrincess:

yes 3/8?

Florisalreadytaken:

No. there are \(5+3=\Large 8\) people in total That means that the denominator is going to be 8 There are 5 seniors Having that info we can write that the chance for the 1st person who is about to get picked, to be a senior \( P(1)=\frac{5}{8}\) The chance for the 2nd person to be a senior is \( P(2)=\frac{5-1}{8-1}\) The odds for the 3rd person to be a senior are \( P(3)=\frac{5-1-1}{8–1-1}\) the probability for all 3 of them can be found by using: \[ P_T=P(1) \cdot P(2) \cdot P(3) \]

Florisalreadytaken:

So we have: \( P(1)=\frac{5}{8}\) \( P(2)=\frac{4}{7}\) \( P(3)= \frac{3}{6} \) \[ P_T=P(1) \cdot P(2) \cdot P(3) \Rightarrow \frac{5}{8} \cdot \frac{ 4}{7} \cdot \frac{3}{6} \] Find the LMP of the 3 denominators

Florisalreadytaken:

Id suggest to just look at the options we’re given. Its not a) because the denominator is an astronomical number b) 8, 7, and 6 do indeed work with 28 c) yeah nah as you can see its a nono. That said, which option is right?